### v1

# Training Small Reasoning LLMs with Cognitive Preference Alignment

Abstract
The reasoning capabilities of large language models (LLMs), such as OpenAI’s o1 and DeepSeek-R1, have seen substantial advancements through deep thinking. However, these enhancements come with significant resource demands, underscoring the need to explore strategies to train effective reasoning LLMs with far fewer parameters. A critical challenge is that smaller models have different capacities and cognitive trajectories than their larger counterparts. Hence, direct distillation of chain-of-thought (CoT) results from large LLMs to smaller ones can be sometimes ineffective and requires a huge amount of annotated data. In this paper, we introduce a novel framework called Critique-Rethink-Verify (CRV), designed for training smaller yet powerful reasoning LLMs. Our CRV framework consists of multiple LLM agents, each specializing in unique abilities: (i) critiquing the CoTs according to the cognitive capabilities of smaller models, (ii) rethinking and refining these CoTs based on the critiques, and (iii) verifying the correctness of the refined results. We further propose the cognitive preference optimization (CogPO) algorithm to enhance the reasoning abilities of smaller models by aligning thoughts of these models with their cognitive capacities. Comprehensive evaluations on challenging reasoning benchmarks demonstrate the efficacy of CRV and CogPO, which outperforms other training methods by a large margin.  [^1] Source codes, datasets and models will be released upon paper acceptance.
Training Small Reasoning LLMs with Cognitive Preference Alignment
Wenrui Cai1,2, Chengyu Wang2thanks:    Corresponding author., Junbing Yan2, Jun Huang2, Xiangzhong Fang1 1 Shanghai Jiao Tong University, Shanghai, China 2 Alibaba Cloud Computing, Hangzhou, China chengyu.wcy@alibaba-inc.com
1 Introduction
The remarkable progress in large language models (LLMs) has revolutionized NLP Zhao et al. (2023). More recently, leading models such as OpenAI’s o1222https://openai.com/o1/ and DeepSeek-R1 DeepSeek-AI (2025) leverage slow thinking mechanisms to solve complex tasks. Despite these capabilities, the scale of these models results in substantial computational demands. Consequently, there is a growing need to facilitate effective reasoning abilities with considerably fewer model parameters.
Figure 1: A motivation example. Large models (right) apply vector-based algebraic abstraction to solve the problem, while small models (left) employ simple formulaic geometric decomposition. This trajectory mismatch underscores the inefficacy of direct CoT distillation across models with substantial capacity gaps.
A straightforward approach to address this challenge is the direct distillation of Chain-of-Thought (CoT) outputs Wei et al. (2022a) or other deep thoughts (such as Tree-of-Thought Yao et al. (2023b)) from larger LLMs to smaller models. This technique is widely applied to improve the capacities of smaller LLMs Hsieh et al. (2023); Shridhar et al. (2022); Li et al. (2023); Yue et al. (2024). However, smaller LLMs [^3] In this work, we regard smaller LLMs as decoder-only language models typically with fewer than 10B parameters. inherently possess different capacities and cognitive trajectories when solving problems compared to their larger counterparts, as illustrated in Figure 1. Similar findings have also been presented in Li et al. (2022); Zhang et al. (2024); Hu et al. (2024); Li et al. (2024). This phenomenon indicates that straightforward distillation of CoTs from larger models can sometimes be ineffective and requires a huge amount of data due to the large capacity gap. Thus, a natural question arises: How can we improve the reasoning abilities of smaller LLMs to fit their own cognitive capacity?
In this paper, we introduce “Critique-Rethink-Verify” (CRV), a novel approach to enhance the reasoning capabilities of smaller LLMs. CRV leverages multiple LLM agents, each with specialized functions and working in synergy. These functions include (i) critique the CoT output by considering the cognitive limits of smaller LLMs, (ii) rethink and refine these CoTs, integrating the feedback received from the previous module, and (iii) verify the accuracy and validity of the refined reasoning paths. Extending the direct preference optimization (DPO) technique Rafailov et al. (2023), we further propose the cognitive preference optimization (CogPO) algorithm to align reasoning abilities with the unique cognitive strengths of smaller LLMs. Ultimately, the reasoning performance of smaller LLMs can be significantly improved.
In the experiments, the effectiveness of our approach is evaluated on several challenging reasoning benchmarks that are difficult for models with limited parameter sizess such as AIME 2024, MATH-500 Lightman et al. (2023), GPQA-Diamond Rein et al. (2023), and LiveCodeBench. The results show that the smaller LLMs trained using the CRV framework and the CogPO algorithm achieve better reasoning performance, outperforming other popular training methods. In summary, we make the following major contributions:
  • We present the CRV framework for training smaller yet powerful reasoning LLMs, leveraging multiple LLM agents where each agent specializes in unique tasks.
  • We propose the CogPO algorithm that enhances the reasoning abilities of smaller models by aligning their reasoning processes with their cognitive capacities.
  • Our evaluations on several challenging reasoning benchmarks demonstrate that combining CRV and CogPO significantly improves the reasoning performance of smaller models.
2 Related Work
2.1 Prompting LLMs to Reason
Prompting strategies to improve reasoning in LLMs have become a critical focus. Initial studies showed that LLMs could perform basic reasoning tasks using meticulously crafted prompts, such as linguistic analysis Chen et al. (2021) and commonsense inference Latcinnik and Berant (2020); Shwartz et al. (2020). These single-step reasoning approaches laid the groundwork for advanced prompting techniques. To name a few, Chain-of-Thought (CoT) Wei et al. (2022b) prompting explicitly guides LLMs through step-by-step reasoning, enabling them to decompose complex problems into manageable intermediate reasoning steps. Tree-of-Thought (ToT) Yao et al. (2023a) prompting introduces a hierarchical structure to reasoning trajectories, allowing models to explore multiple solution paths. Furthermore, self-refine Shinn et al. (2023); Madaan et al. (2023) prompting incorporates verification checkpoints, where models validate intermediate results before advancing.
Figure 2: Overview of our proposed approach. The framework consists of two synergistic phases: (1) Generation of cognitively aligned data using the CRV system for direct distillation using SFT, and (2) CogPO training with dynamic $\beta$ adjustment, considering different levels of mini-tasks. Disclaimer: We use the Qwen logo as our backbones; however, any LLMs with sufficient capabilities can serve as the agents as well.
2.2 Reasoning LLMs
With the advancement of LLMs, model capabilities have steadily improved Chen and Varoquaux (2024); Bansal et al. (2024), even as their scale has been reduced. Models with approximately 7B to 14B parameters show remarkable performance, and their fine-tuning costs have become increasingly feasible. This has led to the emergence of specialized small models tailored for mathematical and code-related reasoning tasks such as Qwen-Math [^4] https://qwenlm.github.io/blog/qwen2.5-math/, Qwen-Coder [^5] https://qwenlm.github.io/blog/qwen2.5-coder-family/, and Macro-o1 Zhao et al. (2024).
Recent studies Shridhar et al. (2023); Yan et al. (2023); Liang et al. (2024); Yuan et al. (2024) have investigated fine-tuning methods to enhance the reasoning abilities of smaller models. By utilizing intermediate reasoning steps, LLMs can iteratively refine their outputs Jiang et al. (2024); Wang et al. (2024); Chen et al. (2025). This progressive methodology enables the development of compact models with reasoning capabilities that rival or exceed those of larger LLMs (over 100B), leading to the release of stronger reasoning models such as QwQ-32B [^6] https://qwenlm.github.io/blog/qwq-32b/ and DeepSeek-R1 DeepSeek-AI (2025). Our work studies the cognitive capacities of small models, beneficial for the community to train more powerful reasoning LLMs.
2.3 Alignment Training
To effectively train LLMs, a reinforcement learning stage is typically employed after the supervised fine-tuning (SFT) phase, which serves to improve the model’s alignment towards certain objectives. Reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has shown effectiveness in aligning LLMs with human feedback. A potential drawback of RLHF is the explicit need for a reward model and the unstable RL training process. Direct preference optimization (DPO) Rafailov et al. (2023) trains LLMs based on chosen and rejected responses. Since the introduction of DPO, several approaches have been proposed to enhance its efficacy and efficiency. For example, CPO Xu et al. (2024) extends DPO to avoid generating adequate but not perfect machine translations. SimPO Meng et al. (2024) simplifies DPO by eliminating the reference model. KTO Ethayarajh et al. (2024) and NCA Chen et al. (2024) develop novel optimization goals that leverage unpaired data for model alignment. Furthermore, SPPO Wu et al. (2024b) employs on-policy sampling to generate preference data, outperforming off-policy DPO methods. In our work, we extend DPO to align reasoning abilities with the cognitive strengths of smaller LLMs.
3 Proposed Approach
We present the techniques of our CRV system and the CogPO training algorithm in detail.
3.1 Overall Framework
Our framework consists of two synergistic phases: (1) SFT with cognitively aligned data using the CRV system, and (2) CogPO with dynamic $\beta$ adjustment. As illustrated in Figure 2, the CRV system first refines training data tailored to the cognitive capacity of smaller LLMs, while CogPO further aligns reasoning preferences through difficulty-aware optimization. This design ensures the model initially acquires capacity-matched reasoning patterns, followed by the refinement of its decision boundaries through gap-sensitive learning.
3.2 The CRV System
The CRV system employs three LLM agents to construct the SFT dataset aligned with the cognitive limits of smaller models. The input to the system is an initial training set $\mathcal{D}_{\text{SFT}}=\{(x,y,r_{\text{orig}})\}$, where the three elements denote the problem, the answer, and the original reasoning process generated by any large reasoning model (e.g., DeepSeek-R1).
3.2.1 Critic
A LLM agent first evaluates the appropriateness of reasoning processes for the target small LLM (denoted as $\pi_{\text{base}}$). For each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{SFT}}$, the Critic assesses $r_{\text{orig}}$ using the criteria of Cognitive Matching Degree, where the Critic checks whether the complexity of $r_{\text{orig}}$ aligns with the reasoning capacity of $\pi_{\text{base}}$. Specifically, the Critic classifies the reasoning processes into three subsets: i) $\mathcal{D}_{\text{easy}}$: cases where the reasoning process is overly simple, failing to guide $\pi_{\text{base}}$ in solving $x$; ii) $\mathcal{D}_{\text{med}}$: cases with adequate steps that enable successful problem solving; and iii) $\mathcal{D}_{\text{hard}}$: cases with redundant or excessively complex reasoning steps that exceed the comprehension of $\pi_{\text{base}}$.
Remarks. An intuitive approach is to use $\pi_{\text{base}}$ as the Critic. However, due to the small parameter size (e.g., 7B), certain CoTs exceeded $\pi_{\text{base}}$’s comprehension, rendering it incapable of reliable difficulty classification. Thus, we leverage the same LLM for the Rethinker and the Verifier (denoted as $\pi_{\text{large}}$) to serve as the Critic, while forcing it to “think” from the perspective of the small model $\pi_{\text{base}}$. Detailed analysis on the Critic choices and prompt templates are provided in experiments and appendix.
3.2.2 Rethinker
A LLM agent $\pi_{\text{large}}$ is tasked with rewriting reasoning processes to achieve the cognitive alignment. For each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{easy}}$, the Rethinker expands $r_{\text{orig}}$ by adding missing logical steps while maintaining the correctness of $r_{\text{orig}}$, i.e.,
$r_{\text{easy*}}=\pi_{\text{large}}(x,y,r_{\text{orig}}).$ (1)
Similarly, for each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{hard}}$, the Rethinker simplifies $r_{\text{orig}}$ by removing redundancies or using simpler method to solve the problem grounded in the correct answer:
$r_{\text{hard*}}=\pi_{\text{large}}(x,y,r_{\text{orig}}).$ (2)
Cases of the rewriting outcomes are shown in Table 8 and Table 9.
3.2.3 Verifier
Finally, we leverage a LLM agent $\pi_{\text{large}}$ to validate the correctness and the rewritten reasoning processes, in order to preserve the high quality of the dataset. It predicts whether $\pi_{\text{base}}$ can derive the correct answer $y$ from the rewritten thoughts for $*r_{\text{easy*}}$ or $*r_{\text{hard*}}$ produced by the Rethinker.
After the verification, incorrect cases are sent back to the Rethinker to be continuously rewritten until they pass verification. In the implementation, a maximum of three iterations are allowed; cases that fail to pass verification after three iterations are discarded. As for the correct cases, we invoke the Critic to make the judgment again (please refer to Figure 2 for the algorithmic flow). The final SFT dataset is composed of verified data: $*\mathcal{D}_{\text{SFT*}}=\mathcal{D}_{\text{med}}\cup\mathcal{D}_{\text{easy*}}\cup\mathcal{D}_{\text{hard*}}$, where $*\mathcal{D}_{\text{easy*}}$ and $*\mathcal{D}_{\text{hard*}}$ represent the rewritten versions of $\mathcal{D}_{\text{easy}}$ and $\mathcal{D}_{\text{hard}}$ that have passed verification and have been re-rated as medium by the Critic, respectively. $*\mathcal{D}_{\text{SFT*}}$ serves as the training set for the first stage of CRV distillation training via SFT.
3.3 Cognitive Preference Optimization
The CogPO algorithm aligns the thought processes of smaller LLMs with their inherent cognitive capacities, following the direct SFT distillation of CoT results refined with the CRV framework.
3.3.1 Preliminaries
Briefly speaking, CogPO is extended from DPO Rafailov et al. (2023) and its variants. Let $y_{w}$ and $y_{l}$ be the chosen and rejected responses for an instruction $x$ (not restricted to reasoning problems addressed in this work), respectively. We further denote $\pi_{\theta}$ as the model to be optimized after SFT and $\pi_{\text{ref}}$ as the reference model. DPO seeks to maximize the following margin:
$M_{\beta}(x,y_{w},y_{l})=\beta\cdot\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ (3)
where $\beta$ is a temperature hyperparameter. Based on $M_{\beta}(x,y_{w},y_{l})$, the DPO loss is defined as:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log\sigma(M_{\beta}(x,y_{w},y_{l})).$ (4)
The settings of $\beta$ are critical to the performance of DPO. $\beta$-DPO Wu et al. (2024a) further adjusts $\beta$ according to $M_{\beta}(x,y_{w},y_{l})$ either at the instance level or batch level, allowing the model to adapt $\beta$ based on the reward differential of the input data.
3.3.2 Algorithmic Description
As noted, DPO and $\beta$-DPO do not require any prior knowledge of how the model learns the user’s preferences. We suggest that this type of prior knowledge is critical for training better smaller reasoning models, as the cognitive trajectories of large and small models often differ Li et al. (2022); Zhang et al. (2024); Hu et al. (2024), which may not be directly reflected in the reward differential. We propose CogPO to align reasoning preferences by encoding more prior knowledge, continuously training on a series of mini-tasks.
Recall that it is probable for the Rethinker in CRV to generate incorrect reasoning processes when asked to rewrite the original thought $r_{\text{orig}}$. The incorrect thoughts are denoted as $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$, based on their origins from $\mathcal{D}_{\text{med}}$, $\mathcal{D}_{\text{easy}}$ and $\mathcal{D}_{\text{hard}}$. These thoughts contain factual errors or invalid reasoning steps, which can mislead $\pi_{\text{base}}$, rendering it unable to solve $x$. Thus, we categorize the properties of all the thoughts we have collected into the following three types: i) $r_{\text{med}}$, $*r_{\text{easy*}}$, and $*r_{\text{hard*}}$: medium-level reasoning processes that are both correct and cognitively suitable for $\pi_{\text{base}}$; ii) $r_{\text{easy}}$ and $r_{\text{hard}}$: easy or hard thoughts that are correct but unsuitable for $\pi_{\text{base}}$; iii) $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$: incorrect reasoning processes with logical flaws or invalid reasoning steps (regardless of the difficulty levels).
Figure 3: An illustration of CogPO, showing the different preference gaps between CoT process pairs and the corresponding mini-tasks.
To define the mini-tasks used for CogPO training, we consider the preference gaps in these three types of thought pairs as follows:
  • Small Gap Mini-task: The thought pairs are from ($*r_{\text{easy*}}$, $r_{\text{easy}}$) and ($*r_{\text{hard*}}$, $r_{\text{hard}}$). Both are correct but differ in complexity (suitable vs. unsuitable for $\pi_{\text{base}}$). We treat $*r_{\text{easy*}}$ and $*r_{\text{hard*}}$ as chosen reasoning processes ($r_{w}$), and $r_{\text{easy}}$ and $r_{\text{hard}}$ as rejected ($r_{l}$).
  • Medium Gap Mini-task: The thought pairs are from ($r_{\text{easy}}$, $\tilde{r}_{\text{easy}}$) and ($r_{\text{hard}}$, $\tilde{r}_{\text{hard}}$). The former are correct but unsuitable, while the latter are completely incorrect. As correctness is more important than suitability for our model, the preference gap of this mini-task should be higher than that in the previous case. For this mini-task, $r_{\text{easy}}$ and $r_{\text{hard}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
  • Large Gap Mini-task: The thought pairs are from ($r_{\text{med}}$, $\tilde{r}_{\text{med}}$), ($*r_{\text{easy*}}$, $\tilde{r}_{\text{easy}}$) and ($*r_{\text{hard*}}$, $\tilde{r}_{\text{hard}}$). Intuitively, the preference gaps should be the largest between suitable and correct thoughts, and incorrect ones. Here, $r_{\text{med}}$, $*r_{\text{easy*}}$ and $*r_{\text{hard*}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
Following our modeling framework, CogPO can be implemented either at the instance level or batch level. For the instance level, each training instance $(x,r_{w},r_{l})$ receives its specific $\beta$ value, as illustrated in Figure 3. The CogPO objective function aggregates these preferences:
$\mathcal{L}_{\text{CogPO}}=-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}}\log\sigma(M_{\beta_{\text{CogPO}}}(x,r_{w},r_{l}))$ (5)
where $\beta_{\text{CogPO}}$ is selected from $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$, depending on the specific types of mini-tasks (with $\beta_{\text{S}}<\beta_{\text{M}}<\beta_{\text{L}}$, corresponding to the three gaps).
As for the batch level, we require our trainer to sample a mini-task for each batch to optimize. The objective function of CogPO is then rewritten as:
$\mathcal{L}_{\text{CogPO}}=$ $-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{S}}}\log\sigma(M_{\beta_{\text{S}}}(x,r_{w},r_{l}))$
$-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{M}}}\log\sigma(M_{\beta_{\text{M}}}(x,r_{w},r_{l}))$
$-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{L}}}\log\sigma(M_{\beta_{\text{L}}}(x,r_{w},r_{l})).$
where $\mathcal{D}_{\text{S}}$, $\mathcal{D}_{\text{M}}$ and $\mathcal{D}_{\text{L}}$ denote the subsets.
Overall, our CogPO algorithm enables granular preference learning: strong regularization ($\beta_{\text{L}}$) for validity discrimination, moderate guidance ($\beta_{\text{M}}$) for suitability alignment, and subtle refinement ($\beta_{\text{S}}$) for reasoning style adaptation. This design provides more control over the thought alignment process, leading to further improvements compared to vanilla SFT distillation or SFT distillation with rewriting based on the CRV system.
Remarks. CogPO can be naturally combined with $\beta$-DPO Wu et al. (2024a). We can redefine the $\beta$ values $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$. Take our instance-level implementation as an example. We define the new instance-level value $\beta_{i}^{*}$ as follows: $\beta_{i}^{*}=\beta_{i}+\alpha\cdot(M_{i}-M_{0})\cdot\beta_{i}$ where $\beta_{i}$ is chosen from $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$ based on the corresponding gap type, $M_{i}$ is the instance-level reward differential, and $M_{0}$ is a predefined threshold as in Wu et al. (2024a). [^7] In our experiment, the combination does not bring substantial improvements, as prior knowledge is more important for our task. Hence, we stick to the usage of $\mathcal{L}_{\text{CogPO}}$.
Dataset/Model Zero-shot SFT CRV+SFT CRV+SFT DPO $\beta$-DPO SimPO CogPO
AIME2024 10.0 20.0 26.7 26.7 23.3 23.3 26.7 30.0
MATH-500 73.6 80.0 83.4 84.0 83.4 83.8 84.2 84.4
GSM8K 89.5 92.3 92.7 91.9 92.6 93.0 92.6 93.3
GPQA Diamond 33.3 37.4 40.9 38.4 40.0 37.4 40.9 40.9
LiveCodeBench V2 30.7 31.3 34.3 34.4 34.4 35.8 36.2 36.6
MMLU 71.9 76.1 76.4 76.5 76.1 76.0 76.5 76.2
OlympiadBench (math-en) 40.1 43.6 45.7 45.8 45.7 46.5 46.0 46.6
Table 1: Performance comparison of various training methods on challenging reasoning benchmarks. The LLM backbone is Qwen2.5-7B-Instruct. The training set is Bespoke-Stratos-17k. Results are shown for zero-shot (without further training), vanilla SFT, CRV (a variant of our approach without thought rewriting), CRV (full implementation), DPO, $\beta$-DPO, SimPO, CogPO methods. DPO, $\beta$-DPO, SimPO and CogPO are conducted on model checkpoints after CRV, using the same preference pair dataset. The metrics are scores of these tasks, with the best results for each dataset in each group marked in bold and the second best underlined.
4 Experiments
To evaluate the effectiveness of the CRV framework and the CogPO algorithm, we conduct a series of experiments on challenging reasoning benchmarks.
4.1 Datasets
In our experiments, we evaluate our work on several benchmarks, including AIME2024888https://maa.org/math-competitions/american-invitational-mathematics-examination-aime, MATH-500 Lightman et al. (2023), GSM8K Cobbe et al. (2021), GPQA Diamond Rein et al. (2023), LiveCodeBench V2 Jain et al. (2024), MMLU Hendrycks et al. (2021), and OlympiadBench (math-en) He et al. (2024). Among these, LiveCodeBench V2 provides a comprehensive evaluation for code generation, MMLU is a large multi-task question-answering benchmark, while the others are mathematical problem sets of varying levels. Since some datasets (e.g., GSM8K) contain their own training sets without deep CoTs, for a unified evaluation, we do not use any training sets from these datasets and instead report the results on the testing sets, regarded as out-of-domain testing. The statistics are shown in the appendix.
For the construction of our training set $*\mathcal{D}_{\text{SFT*}}$, we leverage Bespoke-Stratos-17k [^9] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k, which contains 17K tuples of questions, reasoning processes, and answers directly distilled from DeepSeek-R1 DeepSeek-AI (2025). We also utilize two recently released CoT datasets to conduct supplementary experiments with details presented in the appendix. The first one is Sky-T1-data-17k [^10] https://github.com/NovaSky-AI/SkyThought, which is distilled from QwQ-32B-Preview, where the reasoning abilities are reported to be weaker than those of Bespoke-Stratos-17k. The second one is OpenThoughts-114k [^11] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k, which is distilled from DeepSeek-R1 and verified using a data curation recipe. We have chosen not to use some previously released CoT datasets (e.g., OpenLongCoT-SFT [^12] https://huggingface.co/datasets/SimpleBerry/OpenLongCoT-SFT) due to their significantly weaker reasoning abilities, while some of our benchmarks (e.g., AIME2024, OlympiadBench) are extremely challenging, even for human experts.
Dataset/Model LLaMA-O1 Macro-o1 Bespoke-Stratos-7B Ours OpenThinker-7B Ours
Training Set Size 332K 60K 17K 17K 114K 114K
AIME2024 3.3 6.7 20.0 30.0 31.3 43.3
MATH-500 28.6 38.4 82.0 84.4 83.0 88.4
GPQA Diamond 26.3 31.8 37.8 40.9 42.4 42.9
LiveCodeBench V2 1.57 24.9 36.1 36.6 39.9 46.4
Table 2: Comparison between our model and other small reasoning models in the open-source community. Specifically, we train two versions using our approach on Bespoke-Stratos-17k and OpenThoughts-114k, respectively, where the training sets are in line with Bespoke-Stratos-7B and OpenThinker-7B.
Model Backbone (The Critic) AIME2024 MATH-500 GPQA GSM8K LCB OlympiadBench
Qwen2.5-7B-Instruct 13.3 80.2 40.9 92.7 30.5 43.9
Qwen2.5-32B-Instruct 23.3 82.2 39.9 92.9 33.3 45.1
Qwen2.5-72B-Instruct 20.0 81.8 36.4 93.0 30.5 42.0
DeepSeek-R1-Distill-Qwen-32B 26.7 83.4 40.9 92.7 34.3 45.7
Table 3: Comparison using different backbones as the Critic. GPQA: GPQA Diamond. LCB: LiveCodeBench V2. All the results are produced based on the SFT distillation using rewritten training sets without CogPO training.
4.2 Experimental Details
[…]
4.3 Main Experimental Results and Ablations
Table 1 presents the results of our CRV framework and the CogPO algorithm on various reasoning benchmarks. The training set is Bespoke-Stratos-17k. These results reveal that our CRV framework and the CogPO algorithm significantly enhance the reasoning capabilities of smaller models over traditional methods. In particular, CogPO shows the strongest performance across almost all datasets, indicating its ability to align model thinking processes with their cognitive capacities.
For the ablation studies, we demonstrate that the full implementation of CRV yields gains over its variant CRV, which does not include thought rewriting, in most datasets. This proves the significance of our CRV framework design. However, in a few cases, the performance experiences a slight decline, likely due to the fact that the LLM for the Rethinker inevitably introduces some noise or errors into the training set. Nonetheless, the performance is still higher than vanilla SFT. For alignment training, other algorithms, such as DPO, $\beta$-DPO and SimPO do not necessarily have positive effects on the final scores, illustrating that our CogPO is more suitable training for small reasoning LLMs.
4.4 Comparison Against Other Models
We compare our trained 7B model with other models released in the open-source community. The DeepSeek team DeepSeek-AI (2025) releases distilled models based on a (significantly larger) 800K proprietary SFT dataset, which is not publicly available. To ensure fair comparisons, we consider two reasoning LLMs available before the launch of DeepSeek-R1, namely Macro-o1 Zhao et al. (2024) and LLaMA-O1131313https://github.com/SimpleBerry/LLaMA-O1. We also compare other models trained on datasets distilled from DeepSeek-R1 in the open-source community, including Bespoke-Stratos-7B [^14] https://huggingface.co/bespokelabs/Bespoke-Stratos-7B and OpenThinker-7B [^15] https://huggingface.co/open-thoughts/OpenThinker-7B. The results, along with the sizes of training sets, are shown in Table 2. The results indicate that models trained on datasets distilled from DeepSeek-R1 perform significantly better compared to others.
Using our approach, we train two models using the Bespoke-Stratos-17k and OpenThoughts-114k training sets, respectively. Thus, it is fair to compare our method against that of Bespoke-Stratos-7B and OpenThinker-7B. As seen, our resulting models demonstrate significantly better performance compared to baselines in all benchmarks, which clearly proves the effectiveness of our approach. By leveraging larger training sets, we can surely obtain stronger models based on our approach, which will be left as future work.
4.5 Study on Choices of the Critic
In the previous section, we have claimed that using the small target LLM $\pi_{\text{base}}$ as the Critic does not necessarily produce satisfactory results due to the limited parameter size. In contrast, larger LLMs $\pi_{\text{large}}$ can better “think” like small models. The results of CRV using different backbones as the Critic are shown in Table 3, with the backbones for the Rethinker and the Verifier unchanged. From the results, we can see that they confirm our findings, where larger models consistently perform better than the 7B model in most tasks. Among the three large agents, DeepSeek-R1-Distill-Qwen-32B exhibits the best performance based on majority voting across all the testing sets. Note that we do not leverage the QwQ model series as the Rethinker (instead we use the original Qwen instruct models) because it sometimes generates recursive reasoning loops, not suitable to serve our purpose.
4.6 Study on Model Scales
To study the effectiveness of different parameter sizes on the student models, we further report the performance of Qwen2.5-3B-Instruct and Qwen2.5-14B-Instruct. The experimental settings are identical to those of Qwen2.5-7B-Instruct. The results are presented in Figure 4. We observe that our method is also effective across different model scales. An interesting observation is that the improvement is more significant in Qwen2.5-14B-Instruct compared to Qwen2.5-3B-Instruct. This is because even when we leverage the CRV system to rewrite the CoTs, the large capacity gap between the teacher and student models makes it more challenging for Qwen2.5-3B-Instruct to capture the CoTs by SFT. This finding is also consistent with the recently discovered “distillation scaling law” Busbridge et al. (2025).
(a) Qwen2.5-3B-Instruct
(b) Qwen2.5-14B-Instruct
Figure 4: Experimental results of different sizes of Qwen2.5 models on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
4.7 Study on Other Model Backbones
To evaluate the universality of the proposed approach, we perform additional experiments on multiple backbone models beyond the Qwen2.5 series on Bespoke-Stratos-17k dataset. Figure 5 demonstrates that, for both LLaMA and Mistral models, our approach achieves notable performance gains over the direct SFT baseline across diverse mathematical and coding tasks. These results indicate that the CRV+CogPO framework enables seamless adaptation to different LLM backbones, demonstrating the universality of our approach on various LLM backbones, which also shows the potential of our work to produce stronger models based on other LLMs.
(a) Llama3.1-8B-Instruct
(b) Mistral-7B-V0.3
Figure 5: Experimental results of different model series (Llama3.1-8B-Instruct, Mistral-7B-V0.3) other than Qwen2.5 on AIME2024, MATH500, GPQA Diamond, and LiveCodeBench V2.
Figure 6: Impact of different $\beta$ on AIME2024, GPQA Diamond, LiveCodeBench V2 and OlympiadBench.
4.8 Hyper-parameter Analysis
To evaluate the impact of $\beta$ values in CogPO, we perform a series of experiments with varying $\beta$ values to assess the algorithm’s effectiveness. As shown in Figure 6, the highest performance is attained when assigning tailored $\beta$ values to samples based on their specific gaps, which is a core principle of the CogPO algorithm.
4.9 Case Studies
Due to space limitations, case studies of our model outputs are shown in the appendix. They clearly show how our approach can effectively expand or simplify the reasoning processes based on the critic judgment for the training of small models.
5 Conclusion and Future Work
In this paper, we present the CRV framework, a multi-agent system that effectively trains smaller reasoning LLMs. Our approach leverages the strengths of specialized LLM agents to critique, refine, and verify CoT outputs, leading to enhanced reasoning capabilities in smaller models. The CogPO algorithm further aligns model outputs with their inherent cognitive capacities, significantly improving the performance on several challenging reasoning tasks. In the future, we will i) train and release stronger small models using larger CoT datasets; ii) improve the effectiveness of the CRV framework, especially for much smaller models; and iii) investigate our approach for other domain-specific applications, such as medical diagnosis and legal reasoning.
Limitations
While our proposed framework shows promising results in enhancing the reasoning capabilities of smaller LLMs, several limitations still remain. The CRV framework relies heavily on the contributions of larger models refining the CoT output. This dependency may create challenges in situations where access to larger models is restricted, or these larger models generate incorrect results. In addition, although our framework is designed for smaller LLMs, there remains a ceiling on their performance. By nature, smaller models inherently have reduced capacity to encode complex information and handle nuanced reasoning tasks, which may limit their effectiveness in certain scenarios.
Ethical Considerations
Our work is fully methodological; hence, there are no direct ethical issues. However, smaller models trained on data distilled from larger ones might inherit or exacerbate biased outputs, which can still influence outcomes. We suggest that continuous evaluation of trained LLMs based on ethical guidelines is indispensable.
Appendix A Supplementary Experiments and Discussion
A.1 Results on Weaker CoT Dataset
To demonstrate that our approach is truly superior to vanilla SFT over CoT datasets, we conduct an experiment on the Sky-T1 dataset, which is relatively weaker than Bespoke-Stratos-17k due to the choice of the teacher model (i.e., QwQ-32B-Preview) and the data curation pipeline. The results are presented in Table 4. As shown, in some cases, the SFT baseline cannot even beat the zero-shot performance because the 7B backbone is already strong. Hence, training on Sky-T1 with our small LLM backbone may not improve reasoning abilities. This negative finding is also consistent with their own blog regarding the model size and data quality [^16] https://novasky-ai.github.io/posts/sky-t1/. Nonetheless, by comparing our method with the SFT baseline, we can observe clear improvement, which demonstrates that our approach has the efficacy to enhance the reasoning abilities of small models in various scenarios.
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 16.7 20.0
MATH-500 73.6 73.2 77.0
GPQA Diamond 33.3 28.8 36.9
LiveCodeBench V2 30.7 20.9 23.9
Table 4: Performance comparison using Sky-T1-data-17k as the training set.
A.2 Results on Larger CoT Dataset
We further evaluate the performance of our method using OpenThoughts-114k as the training set, which is much larger than other training sets. This dataset is distilled from DeepSeek-R1 and goes through several quality verification steps. The results are presented in Table 5. It can be seen that our method ultimately exhibits exceptionally strong reasoning performance, significantly surpassing SFT on all benchmarks. This underscores the scalability and generalizability of our CRV+CogPO framework to larger datasets.
Dataset/Model Zero-shot SFT Ours
AIME2024 10.0 31.3 43.3
MATH-500 73.6 83.0 88.4
GPQA Diamond 33.3 42.4 42.9
LiveCodeBench V2 30.7 39.9 46.4
Table 5: Performance comparison using OpenThoughts-114k as the training set.
A.3 Design Choice of the Critic
An initial, straightforward method is to employ $\pi_{\text{base}}$ as the Critic. Unfortunately, due to the limited reasoning capabilities inherent to the small model, it consistently faces difficulties in distinguishing the complexity levels effectively in our datasets. Consequently, we utilize a more robust and larger LLM for both the Rethinker and Verifier (referred to as $\pi_{\text{large}}$), which then functions as the Critic. This involves guiding the large model to simulate the cognitive approach of the smaller model, $\pi_{\text{base}}$. The prompt template of the Critic is shown in Table 10. This setting is akin to educational practices, where professors, instead of students, customarily curate academic content across a spectrum of difficulty levels due to their broader knowledge base. As shown in Table 3, the experiments clearly demonstrate the superior evaluative proficiency of the large model, confirming its advantage in categorizing CoT complexity from the perspective of the smaller model efficiently.
Another design choice is that, we divide the CoT processes into three categories: “easy”, “medium” and “hard”, instead of two categories: “correct” and “incorrect”. This is because for incorrect CoT processes, this binary classification fails to distinguish the root cause of errors (i.e., whether the CoT is overly simplified or overly complex), which is critical for determining appropriate refinement strategies (e.g., expansion for overly simplified processes vs. simplification for overly complex ones). Determining the correctness of CoT processes should be the responsibility of the Verifier.
Appendix B Testing Set Statistics
The sizes of our testing sets are summarized in Table 6.
Dataset Size
AIME2024 30
MATH-500 500
GSM8K 1319
GPQA Diamond 198
LiveCodeBench V2 511
MMLU 14042
OlympiadBench (math-en) 674
Table 6: Testing set statistics.
Appendix C Hyperparameter Settings and Training Time
The default hyperparameters for model training and baselines are shown in Table 7. On the Bespoke-Stratos-17k dataset, for the 3B model, we use a single node with 8 A800 GPUs (80GB), with an approximate training time of 4 hours. For the 7B model, we use a single node with 8 A800 GPUs (80GB), with a training time of about 5 hours. For the 14B model, we use 4 nodes, each with 8 A800 GPUs (80G), resulting in a training time of approximately 14 hours.
Hyperparameter Value
CRV Stage
Batch size 96
Learning rate 1e-5
Learning epoch 3.0
CogPO Stage
Batch size 96
Learning rate 5e-7
Learning epoch 1.0
SFT (Baseline)
Batch size 96
Learning rate 1e-5
Learning epoch 3.0
DPO (Baseline)
Batch size 96
Learning rate 5e-7
Learning epoch 1.0
$\beta$ 0.1
SimPO (Baseline)
Batch size 96
Learning rate 5e-7
Learning epoch 1.0
$\beta$ 2.0
$\gamma$ 0.3
Table 7: Training hyperparameters.
Problem Find the inverse of matrix $A=\begin{bmatrix}2&1\\ 1&2\end{bmatrix}$
Answer $A^{-1}=\frac{1}{3}\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$
Original reasoning process Calculate determinant $\det(A)=3$, thus $A^{-1}=\frac{1}{3}\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$
(correct but unsuitable)
Extended reasoning process Compute determinant: $2\times 2-1\times 1=3$
(correct and suitable) Construct adjugate: $\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$
Normalize: $A^{-1}=\frac{1}{3}\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$
Incorrect reasoning process Swap diagonal elements: $\begin{bmatrix}2&1\\ 1&2\end{bmatrix}\rightarrow\begin{bmatrix}2&-1\\ -1&2\end{bmatrix}$
Table 8: Case study of how the reasoning process is extended.
Problem Find the area of a triangle with vertices at $(0,0)$, $(3,0)$, and $(0,4)$
Answer $6$
Original reasoning process Vector Representation: $\vec{AB}=\begin{pmatrix}3\\ 0\end{pmatrix},\vec{AC}=\begin{pmatrix}0\\ 4\end{pmatrix}$
(correct but unsuitable) Determinant Method: $\text{Area}=\frac{1}{2}\left|\begin{vmatrix}3&0\\ 0&4\end{vmatrix}\right|=\frac{1}{2}(12)=6$
Simplified reasoning process Recognize right-angled triangle: $\text{Base}=3,\ \text{Height}=4$
(correct and suitable) Apply elementary formula: $\text{Area}=\frac{1}{2}\times\text{Base}\times\text{Height}=\frac{1}{2}\times 3\times 4=6$
Incorrect reasoning process Area=12(Sum of $\text{Area}=\frac{1}{2}(\text{Sum of sides})=\frac{1}{2}(3+4+5)=6$
Table 9: Case study of how the reasoning process is simplified.
Role Prompt Template
Critic You are a highly capable evaluator.
Your task is to assess the given reasoning process from the perspective of a small language model (e.g., 7B).
Specifically, determine whether the reasoning process provides sufficient detail for a small model to solve the
problem, or whether it is too simplistic (i.e., lacking critical details) or too complex (i.e., containing
unnecessary or confusing steps).
Difficulty Definitions (from the perspective of a small model):
- Easy: The reasoning process is overly simplistic relative to the problem’s difficulty; it omits essential details that
a small model needs to solve the problem.
- Medium: The reasoning process is appropriately balanced, offering enough detailed guidance.
- Hard: The reasoning process is overly complex, with extraneous or convoluted steps that could hinder a small
model’s ability to follow it.
Output Format:
You must output exactly one word: easy, medium, or hard. Do NOT provide any additional text, explanation.
Rethinker You are a helpful assistant who is highly skilled at extending reasoning processes.
(easy) Given a problem ,its answer and its reasoning process, your task is to extend the reasoning process by adding
necessary details and intermediate steps so that a small language model (e.g., a 7B model) can follow the
extended reasoning process to solve the problem.
If the original reasoning process is divided into multiple steps separated by two newline characters, your output
must preserve this formatting.
You must output ONLY the extended reasoning process with no additional explanation or commentary.
Rethinker You are a helpful assistant who is highly skilled at simplifying reasoning processes.
(hard) Given a problem, its answer and its reasoning process, your task is to simplify the reasoning process so that
a small language model (e.g., a 7B model) can reliably follow the steps to solve the problem.
If the original reasoning process is divided into multiple steps separated by two newline characters, your output
must preserve this formatting.
You must output ONLY the simplified reasoning process with no additional explanation or commentary.
Verifier You are a highly capable Verifier.
Your task is to assess a given reasoning process based on a problem, its answer, and the reasoning process.
Specifically, determine whether the reasoning process is sufficient and accurate for a small model (e.g., 7B)
to reach the correct answer.
If the reasoning process aligns with the model’s cognitive abilities and can correctly guide the small model
to the solution, output YES.
If the reasoning process is flawed or does not match the model’s capabilities such that it fails to guide
the model to the correct answer, output NO.
You must output exactly one word: YES or NO. Do NOT provide any additional text, explanation.
Table 10: Prompt templates for the CRV system.
Appendix D Case Studies
Case studies are presented in Table 8 and Table 9.
Appendix E Prompt Templates
Prompt templates of the Critic, the Rethinker and the Verifier in our CRV system are shown in Table 10.